### v1

# LEAM: A Prompt-only Large Language Model-enabled Antenna Modeling Method

Abstract
Antenna modeling is a time-consuming and complex process, decreasing the speed of antenna analysis and design. In this paper, a large language model (LLM)- enabled antenna modeling method, called LEAM, is presented to address this challenge. LEAM enables automatic antenna model generation based on language descriptions via prompt input, images, descriptions from academic papers, patents, and technical reports (either one or multiple). The effectiveness of LEAM is demonstrated by three examples: a Vivaldi antenna generated from a complete user description, a slotted patch antenna generated from an incomplete user description and the operating frequency, and a monopole slotted antenna generated from images and descriptions scanned from the literature. For all the examples, correct antenna models are generated in a few minutes. The code can be accessed via https://github.com/TaoWu974/LEAM.
Index Terms:
Large language model, antenna, electronic design automation, GPT, antenna modeling, human–computer interaction
I Introduction
In antenna research and applications, there is a demanding need for exploring novel antenna structures and reproducing designs from the literature. Antenna modeling is often unavoidable in this process, costing a lot of time and effort. Depending on the user’s experience, modeling a complex antenna could take from several tens of minutes to hours, and the number of such antenna modeling tasks grows rapidly with the growth of the antenna industry. It would be ideal to reduce or remove this repetitive and error-prone work. This inspires us to automate antenna modeling, speeding up the antenna design and innovation process.
It is routine to use the graphical user interface (GUI) in electromagnetic (EM) modeling and simulation tools for antenna modeling and visualization. These visualized antenna structures can be represented in macro files, which are command lines describing the model. For example, CST Microwave Studio employs Visual Basic for Applications (VBA) macros, and HFSS supports IronPython macros. Hence, by generating appropriate macro files, the antenna modeling process can become automated.
However, macro files are complex, and manual scripting is even more time-consuming than drawing the model using GUI. For example, importing material to an object is a straightforward click using the GUI of EM simulation tools, but tens of parameters are needed in the macro file. The gap between the simple language description of the user and the complex macro file that can correctly define the antenna model becomes the bottleneck, which has not been solved till now. However, a recent innovation, Large Language Models (LLMs), provides a promising solution, bridging simple descriptions in prompt (a prompt is the text input to the LLM) and complex macro files by its code generation [1], geometry understanding [2], parameterization [3] and learning ability.
Although have not been used in antenna modeling, LLMs are investigated for automated macro scripting for 3D mechanical structure modeling. Some pioneer research works and findings are as follows. [4] shows weak prompts can lead to poor responses and task failures. [5] shows that LLMs with vision capabilities (i.e., image recognition) outperform text-only LLMs as the modeling complexity increases. [6] collects macro templates and benchmark test cases to assess LLM-based modeling techniques. SolidWorks employs LLMs in modeling mechanical components [7].
Despite the advancements in applying LLMs to 3D modeling, two limitations are restricting their applications in antenna modeling: (1) There is limited research considering multi-modal input (i.e., not only textual descriptions, but also image input) [5], but is highly needed in LLM-enabled antenna modeling, especially for reproducing models from the literature. (2) Antenna modeling often consists of a series of solids’ modeling and boolean operations among them, which is quite different from mechanical modeling’s process of selecting one model and configuring its parameters [5].
To address these challenges, the LLM-enabled antenna modeling method (LEAM) is proposed. The macro scripting process is divided into systematic steps, and a set of LLM tools based on OpenAI LLM Application Programming Interface (APIs) with designated prompts including responses from other LLM tools are designed for these steps. The input of LEAM is textual descriptions and images (if available), and the output is the antenna model. Currently, CST Microwave Studio is used as the EM simulator. To the best of our knowledge, this research is the first work investigating LLM-enabled antenna automated modeling. Moreover, LEAM is a prompt-only method, avoiding complex and time-consuming fine-tuning (i.e., additional training using specific data for antennas) or retrieval-augmented generation (i.e., providing knowledge databases, such as documents, and books), improving usability.
II The LEAM Methodology
II-A The workflow of LEAM
In LEAM, the antenna modeling approach is divided into systematic steps, and for each step, a tool is created. The tools are shown in Fig. 1 and Table I. The user employs the tools mostly sequentially, except in the first step, there is selection, which will be detailed later.
Figure 1: LEAM workflow
TABLE I: LEAM Tools with I/O, and LLM Model
Tool Input Output LLM
Weak D2S Text, Image (if available) Solid_List.txt o1
Strong D2S Text, Image (if available) Solid_List.txt o1
Parameterize Text, Image (if available) Para.bas o1/4o
Dimension Text, Image (if available) Solids_Dims.txt o1/4o
Materials Solid_List.txt Materials.bas 4o
Model3D Solids_Dims.txt & *.bas 3D_Model.bas o1
Model2D+ Solids_Dims.txt & *.bas 2D+_Model.bas o1
Boolean Solids_Dims.txt & *.bas Boolean.bas o1
UpdatePara *.bas UpdatePara.bas o1
RunVBA *.bas Model.cst N/A
As shown in Fig. 1, the modeling process is divided into four stages, which are identifying solids, defining solids, modeling solids, and executing macros. (In this work, CST Microwave Studio is used, where solid is the fundamental modeling unit.) In the first stage, if the provided description is sufficient, the strong descriptions to solids (Strong D2S) tool will be used; otherwise, the weak descriptions to solids (Weak D2S) tool will be used. Both tools have the same output, which lists all solids with their materials, types (e.g., substrate, feeding line, slot, etc.), geometric shapes (e.g., brick, cylinder, extruded 2D shape, etc.), and positions.
The second stage is to define solids. Using the solid list obtained from the first stage, the Parameterize tool is then employed to define parameters and generate a Para.bas file to import parameters into the solids. After that, with defined parameters and descriptions (with images if available), the solid list is updated by adding detailed dimensions and positions. Next, the Materials tool queries the EM modeling tool’s material library to retrieve the materials specified in the solid list and subsequently generates the Materials.bas to import these materials into the solids.
The third stage is to model solids, which can be divided into 3D shapes or 2D+ shapes. Given the solid list with dimensions, parameters, and materials, the Model3D tool identifies and models the 3D shapes, and produces the 3D_Model.bas file. 2D+ shape refers to extruded or rotated 2D shapes. The Model2D+ tool first models closed 2D shapes and then extrudes and/or rotates these 2D shapes into 3D shapes, and produces the 2D+_Model.bas file. Then, given the modeled solids in the above bas files and the solid list with dimensions, the Boolean tool determines sequential boolean operations. For example, adding the connected solids, subtracting the slot solids.
The last stage is executing macro files. The RunVBA tool executes all bas files obtained before, including Para.bas, Materials.bas, 3D_Model.bas, 2D+_Model.bas, and Boolean.bas. These macros are saved to the history list of CST Microwave Studio. The automated antenna modeling is completed. After this step, an optional tool is the UpdatePara tool, which is used to change parameters. When the user provides new parameter sets, the UpdatePara.bas is generated and executed to update the antenna model.
II-B LLM tool creation
As said above, 9 LLM tools and a RunVBA tool are defined in LEAM. Fig. LABEL:fig:_tool_I/O shows a tool structure using the Dimension tool as an example. Each tool accepts pre-defined input, input from former tools, and user input. The pre-defined input is the major effort to link LLM with antenna modeling, including role and task prompts, functional instructions, and knowledge files. The role and task prompts provide a task-oriented identity to help the LLM align with task requirements [8]. The functional instructions define the task details and output rules (e.g., ”Output VBA code only”) [9]. Knowledge files are examples of macros and formatted output to guide the LLM.
Figure 2: The Dimension tool
It is worth noting that various LLM models are carefully selected for each tool. With the advancement of reasoning LLMs such as OpenAI o1 [10] and Deepseek R1 [11], LLMs can now build more complex workflows using the Chain of Thought (CoT) [10], aligning seamlessly with complex scripting requirements of antenna modeling. Our pilot experiments show that OpenAI o1 is better at understanding a topology and dismantling it into basic geometric shapes, while GPT-4o outperforms OpenAI o1 in image recognition. Hence, to trade off the reasoning ability and image recognition ability, the general principle is that the base model (GPT-4o-2024-11-20) is used for most multi-modal input and the reasoning model (o1-2024-12-17) is used for text-only input. However, when advanced reasoning is critical, such as the D2S tool, OpenAI o1 is employed even though the input is multi-modal.
II-C Demonstration of LEAM by a simple case
To demonstrate the use of LEAM, a rectangular patch with an L-shape slot example (Fig. 3) is used.
Figure 3: Example: L-slot on the patch. $Patch_{L}$ = 8, $Patch_{W}$ = 10, $Slot1_{L}$ = 4, $Slot1_{W}$ = 1, $Slot2_{L}$ = 6, $Slot2_{W}$ = 1, $Slot1_{offset}$ = 2. (unit: mm.)
In the first stage, user descriptions are input. For example, ”Draw an L-shaped 2D polygon in the XY-plane with the desired corner coordinates (e.g., starting at (LX1, LY1), turning at (LX2, LY2), etc.). Close the polygon and extrude in the Z direction.”. The outcome of this stage is a solid list including two solids, which are a brick for the patch and an extruded 2D shape for the slot.
In the second stage, a solid list with dimensions and two macro files are generated: Para.bas to import 7 parameters ($Patch_{W}$, $Patch_{L}$, $Slot1_{W}$, $Slot1_{L}$, $Slot2_{W}$, $Slot2_{L}$, $Slot_{Offset}$) and Materials.bas to import copper (pure) as the patch’s material. The position of the brick for the patch is defined by the coordinate ranges (”Xmin=0, Xmax=PatchW, Ymin=0, Ymax=PatchL, Zmin=0, Zmax=0.035”), while the L-shape slot is defined by six points forming the L-shape polygon whose Z range is of [0, 0.035]. The descriptions are as follows.
Start from [$\mathrm{Slot_{Offset}}$,$\mathrm{Slot_{Offset}}$]
To [$\mathrm{Slot_{Offset}+Slot2_{L}}$, $\mathrm{Slot_{Offset}}$]
To [$\mathrm{Slot_{Offset}+Slot2_{L}}$, $\mathrm{Slot_{Offset}+Slot2_{W}}$]
To [$\mathrm{Slot_{Offset}+Slot1_{W}}$, $\mathrm{Slot_{Offset}+Slot2_{W}}$]
To [$\mathrm{Slot_{Offset}+Slot1_{W}}$, $\mathrm{Slot_{Offset}+Slot1_{L}}$]
To [$\mathrm{Slot_{Offset}}$, $\mathrm{Slot_{Offset}+Slot1_{L}}$]
To [$\mathrm{Slot_{Offset}}$,$\mathrm{Slot_{Offset}}$]
In the third stage, a 3D_Model.bas for the patch, a 2D+_Model.bas for the L-shape slot, and a Boolean.bas for subtracting the L-shape slot from the patch are generated. To model the 2D+ L-shape slot solid, 2D+_Model.bas first creates an L-shape polygon and then extrudes it to a height of 0.035 (the same as the patch). The fourth stage executes the macro files and the resulting antenna model is in Fig. 3.
III Experimental Results
In this section, three case studies are used to demonstrate LEAM. The first case aims to model a Vivaldi antenna with detailed description text input by the user. The second case aims to model a patch antenna operating at 2.45 GHz but the description text input by the user is incomplete. The third case aims to reproduce the model of a monopole antenna using descriptions and images scanned from [12]. For all the test cases, LEAM successfully generated the correct antenna model in a few minutes.
III-A Case 1: The Vivaldi antenna modeling
[13] indicates that the exponential non-uniform slot plays an important role in reducing the antenna’s physical dimension, and the circular stub on the back of the substrate provides better impedance matching for a smooth microstrip-to-slotline transition. Inspired by this, a Vivaldi antenna is proposed with a substrate (30 mm $\times$ 20 mm), a slotline defined by two symmetric splines and a circular slot on the patch above the substrate, and a feeding structure of two connected feedlines with a circular stub under the substrate. The splines are defined by 20 points, whose coordinates are as follows.
$Left$ $:(x_{i},y_{i}),\quad\forall i=1,2,\dots,20$ (1)
$Right$ $:(30-x_{i},y_{i})\quad\forall i=1,2,\dots,20$
[…]
A complete text description of the Vivaldi antenna is prepared. The description for the slot curve of the patch is as follows.
  • Model two symmetric splines and connect their top and bottom points to form a closed shape to extrude.
  • The left-spline is defined by 20 points, whose y axis are 20, 19, … 1, and the X axis are variables. (Here we will have 20 variables).
  • The point’s $x$ axis is ascending within the range (0, 15), exclusive. Meanwhile, the right half’s $x$ value is 30-$x\_i$.
Using the complete text description for the Strong D2S Tool, 30 parameters and 2 materials are imported. 6 3D solids (substrate, front patch, front cylinder, back rectangular#1, back rectangular#2, back cylinder) and a 2D+ solid (slotline) are modeled and shown in Fig. 4, whose parameters are updated by the UpdatePara tool to form the Vivaldi shape. The generated antenna model is shown in Fig. 4.
(a) Front view
(b) Back view
Figure 4: Vivaldi antenna. SL = 30, SW = 20, $r_{1}=2$, $gap_{1}=1$, $L_{1}=7$, $L_{2}=8$, $gap_{2}=5$, $W_{1}=2$, $W_{2}=1$, $r_{1}=2$. (unit: mm.)
III-B Case 2: The patch antenna modeling
Unlike case 1, the description for case 2 is incomplete. The only prompt is ”I want to design a rectangular-patch antenna with a rectangular slot working at 2.45 GHz”. The Weak D2S tool in LEAM identifies five 3D solids (substrate, patch, slot, feedline, and ground plane) with 2 materials and 4 parameters. The modeled antenna is shown in Fig. 5, with its $S_{11}$ response exactly at 2.45 GHz shown in Fig. 6. To identify the appropriate dimensions of the patch, LEAM builds CoT and applies the transmission line model in [14]. This case demonstrates that even without a complete description, LEAM can make use of CoT in LLM technology and help users who are not antenna experts.
Figure 5: Slotted patch antenna @ 2.45GHz. PatchW = 38, PatchL = 28, SlotW = 2, SlotL = 10. (unit: mm.)
Figure 6: Simulated $S_{11}$ response for the slotted patch antenna.
III-C Case 3: The monopole antenna modeling
The third case is a slotted monopole antenna from the literature. Its description, topology figure, and parameter list figure are scanned from [12] (Fig. 7(a)). The strong D2S tool identifies 7 3D solids (substrate, circular patch, horizontal slot, vertical slot, left ground plane, and right ground plane) with 12 parameters (10 parameters in the figure and 2 thickness parameters in the text) and 2 materials. Moreover, we add a prompt to the Parameter tool to correct a typo in [12] ($S_{L}=M_{L}+DP_{R}+0.2$). The modeled antenna is shown in Fig. 7.
(a) Descriptions of the slotted monopole antenna, retrieved from Section IV, second paragraph of [12].
(b) Monopole
Figure 7: Slotted monopole antenna. $DP_{R}$ = 6.58, $S_{W}$ = 13.43, $SL_{T}$ = 1, $SL_{V}$ = 7.9, $SL_{H}$ = 7.9, $M_{L}$ = 25.08, $RP_{L}$ = 6.67, $M_{W}$ = 1.2, $M_{G}$ = 0.3, $S_{L}$ = 31.86, $RP_{W}$ = 5.815. (unit: mm.)
IV Conclusion
LEAM, an open-source, prompt-only LLM-enabled antenna modeling technique designed for automated antenna modeling, has been presented. Using three cases with complete description, weak description, and multi-modal description (i.e., image and text scanned for the literature), the effectiveness of LEAM has been shown. Particularly, LEAM realized prompt-only (fine-tuning and retrieval-augmented generation are not needed). The effectiveness comes from the proposed systematic framework and the defined tools. To the best of our knowledge, LEAM is the first LLM-based antenna modeling tool. Future works include enhancing LEAM’s compatibility with more LLMs (e.g., DeepSeek, Llama) and EM simulation software (e.g., Ansys HFSS, MATLAB Antenna Toolbox).